A ​$1700 certificate of deposit held for 67 days was worth ​$1714.87. What simple interest rate was​ earned? Assume 360 days in a year.
Answer #1

Let us assume that interest rate is R

Interest earned = 1714.87-1700 = $14.87

Interest earned = Principal amount *R* d/360 (d=number of days)

By rearranging the above formula we get,

Interest Rate R = (14.87/1700) *360/67

Interest Rate = 4.7%